# FineServe: Precision-Aware KV Slab and Two-Level Scheduling for Heterogeneous Precision LLM Serving

### v1

@@ 6.3. Local Scheduler: Adaptive Batching
The local scheduler runs on each node and enforces the TTFT SLO for every model while making efficient use of available GPU memory and compute resources. It maintains a separate queue for each model. Inputs include the set of waiting requests, the set of currently running requests, the current time, and per-model hard caps on concurrent requests () and batched prompt tokens (), both derived from current remaining capacity in compute and memory.
Algorithm 2: SLO-aware adaptive batching. The local scheduler processes waiting requests of a model by setting each request’s deadline as its arrival time plus the model’s SLO. Requests that cannot meet their deadline, even if served immediately, are dropped at once. The remaining requests are sorted in increasing deadline order, following EDF policy (Lines 1-4). The request with the earliest deadline anchors the next batch (Line 8). Additional requests are then considered in deadline order and tentatively added as long as the predicted prefill finish time of the combined batch remains within the anchor’s deadline (Lines 9-12). From this tentative batch, requests are admitted only if they also satisfy system constraints on the maximum number of concurrent requests () and the maximum number of batched tokens () (Lines 11). The resulting feasible batch is appended to the tentative schedule, though only the first batch is dispatched immediately (Lines 13-17, 25). If no request can be added without missing the earliest deadline, the scheduler creates slack by removing the longest-running request from the tentative schedule. The batch containing that request is updated by subtracting its original prefill time and recomputing its duration without the request. If the batch becomes empty, it is removed from the schedule altogether. This adjustment frees capacity to rebuild a new tentative batch for the current earliest deadline, ensuring that admitted requests can be served within their SLOs (Lines 18-24).
Design rationale. Our algorithm fundamentally builds on the Moore-Hodgson algorithm (Moore, 1968) to meet the SLO, but adapts it to fit chunked prefill settings. The Moore-Hodgson algorithm is specialized for serving requests individually. As a result, the request sequence it produces is optimal only when the batch size is one. In this case, however, chunked prefill cannot be exploited, leading to performance degradation. When the sequence is executed with a batch size greater than one, the algorithm becomes theoretically sub-optimal because processing time grows with the entire batch, which increases the likelihood of SLO violations. To address this, we modified the algorithm itself so that it forms batches while explicitly ensuring that the SLO of the earliest-deadline request in each batch is not violated.
@@ 7.3. Ablation Study: Two-level Scheduler
Local scheduler evaluation. To isolate the effect of the local scheduler, we evaluate scenarios where a single model is placed on a GPU and the KV cache is managed statically. We compare three schedulers: FineServe, Prism⋆, and FCFS. The experiment uses the Llama-3.1 8B FP16 model under two request rate scales: 1.0 in Figure 13(a) and 1.33 in Figure 13(b). For each case, we set the SLO latency baseline to the P90 TTFT latency attained by FCFS. Results show that FineServe achieves up to 33.57% higher SLO attainment than FCFS and up to 1.63% higher than Prism⋆. Both FineServe and Prism⋆ outperform FCFS due to their Moore-Hodgson-based scheduling, and the slight but consistent advantage of FineServe over Prism⋆ stems from its more efficient use of chunked prefill. Mainly, compared to Prism⋆, the highest difference in SLO attainment occurred in the case of the request rate scale 1.33 and the SLO scale 3.0. Due to the large SLO scale, Prism⋆ tends to batch a large number of requests, and as this batch of requests proceeds with chunked prefill, the likelihood of SLO violations increases. This demonstrates the practical benefit of FineServe ’s adaptive batching in meeting latency SLOs under higher request rates.
@@ 8. Related Work
KV cache memory management. Deployment frameworks based on PagedAttention mechanism, such as vLLM (vLLM, 2025) and SGLang (SGlang, 2025), maintain a private KV block pool for each model. Hence, models sharing a GPU do not borrow memory from one another. Prism (Yu et al., 2025) proposes cross-model sharing by using CUDA VMM to preserve contiguous views while mapping physical memory on demand. Frequent map and unmap calls introduce driver overhead, and heterogeneous block layouts can amplify fragmentation (Prabhu et al., 2025). FineServe differs by pre-allocating a shared KV tensor per GPU and dividing it into uniform slabs indexed by KV block size. Compatible models share a slab pool and suballocation stays in user space with tensor views.

### v2

@@ 6.3. Local Scheduler: Adaptive Batching
The local scheduler runs on each node and enforces the TTFT SLO for every model while making efficient use of available GPU memory and compute resources. It maintains a separate queue for each model. Inputs include the set of waiting requests, the current time, and per-model hard caps on concurrent requests () and batched prompt tokens (), both derived from the current remaining capacity in compute and memory.
Design rationale. Our algorithm fundamentally builds on the Moore-Hodgson algorithm (Moore, 1968) to meet the SLO, but adapts it to fit chunked prefill settings. The Moore-Hodgson algorithm is specialized for serving requests individually. As a result, the request sequence it produces is optimal only when the batch size is one. In this case, however, chunked prefill cannot be exploited, leading to performance degradation. When the sequence is executed with a batch size greater than one, the algorithm becomes theoretically sub-optimal. The reason is as follows: The algorithm tends to form larger batches of requests because it considers the processing time of each individual request, which is shorter than that of the entire batch. This leads to the conclusion that more requests will meet their deadlines. However, with chunked prefill, the processing time increases with the entire batch, which in turn increases the likelihood of SLO violations for requests with earlier deadlines. To address this issue, we modified the algorithm itself so that it forms a batch of requests to be served while explicitly ensuring that the SLO of the earliest-deadline request within the batch remains unviolated when the entire batch is processed by chunked prefill.
Algorithm 2: SLO-aware adaptive batching. The local scheduler processes waiting requests of a model by setting each request’s deadline as its arrival time plus the model’s SLO (line 4). Requests that cannot meet their deadline, even if processed individually and immediately, are dropped at once (lines 7-8). By dropping these requests, GPU memory and compute resources of the system become more available, thereby enhancing the likelihood of meeting the deadlines for other requests. The remaining requests are sorted in ascending order by their deadlines, following the EDF policy, and form a tentative batch of requests to be served (lines 5-6, 9). The request with the earliest deadline acts as the anchor for the tentative batch of requests (line 11). The longest-running request is removed from the tentative batch as long as the predicted prefill finish time of the batch exceeds the anchor’s deadline (lines 12-15).
To fully leverage chunked prefill, we focus on the prefill finish time of the entire batch of requests rather than individual requests, and this adjustment frees capacity to rebuild a new tentative batch, guaranteeing that admitted requests can be served within their SLOs. Finally, from the tentative batch, the batch of requests to be served is formed, ensuring that it adheres to the system constraints by not exceeding the maximum number of concurrent requests () and the maximum number of batched tokens () (line 16). In 7.3, we explain the effectiveness of our algorithm.
@@ 7.3. Ablation Study: Two-level Scheduler
Local scheduler evaluation. To isolate the effect of the local scheduler, we evaluate scenarios where a single model is placed on a GPU and the KV cache is managed statically. We compare three schedulers: FineServe, Prism⋆, and FCFS. The experiment uses the Llama-3.1 8B FP16 model under two request rate scales: 1.66 in Figure 13(a) and 2.0 in Figure 13(b). For each case, we set the SLO latency baseline to the P90 TTFT latency achieved by FCFS. Results show that FineServe achieves up to 45.34% higher SLO attainment than FCFS and up to 8.87% higher than Prism⋆. Both FineServe and Prism⋆ outperform FCFS due to their Moore-Hodgson-based scheduling, and the consistent advantage of FineServe over Prism⋆ stems from its more efficient use of chunked prefill. Mainly, compared to Prism⋆, the highest difference in SLO attainment occurred in the case of the request rate scale 2.0 and the SLO scale 3.0. Due to the large SLO scale, Prism⋆ tends to batch a large number of requests, and as this batch of requests proceeds with chunked prefill, the likelihood of SLO violations increases. This demonstrates the practical benefit of FineServe ’s adaptive batching in meeting latency SLOs under higher request rates.
@@ 8. Related Work
KV cache memory management. Deployment frameworks based on PagedAttention mechanism, such as vLLM (vLLM, 2025) and SGLang (SGlang, 2025), maintain a private KV block pool for each model. Hence, models sharing a GPU do not borrow memory from one another. Prism (Yu et al., 2025) proposes cross-model sharing by using CUDA VMM to preserve contiguous views while mapping physical memory on demand. Frequent map and unmap calls introduce driver overhead, and heterogeneous block layouts can amplify fragmentation (Prabhu et al., 2025). FineServe differs by pre-allocating a shared KV tensor per GPU and dividing it into uniform slabs indexed by KV block size. Compatible models share a slab pool and suballocation stays in user space with tensor views. On top of this, a two-level scheduler places models and sets batching using a memory-efficiency signal so remaining capacity is used where it produces the most tokens while meeting SLOs.